OpenStudy (anonymous):

Find dy/dx given \[y=\sqrt{3x+\sqrt{x^2+1}}\]

zepdrix (zepdrix):

root within a root? :) oo this one looks kind of fun!

zepdrix (zepdrix):

Do you remember your derivative for square root x? It's good to have it memorized instead of converting to rational exponent, it shows up often.

OpenStudy (anonymous):

no

zepdrix (zepdrix):

\[\Large\rm \frac{d}{dx}\sqrt x=\frac{1}{2\sqrt x}\]

zepdrix (zepdrix):

\[\Large\rm y=\sqrt{3x+\sqrt{x^2+1}}\]Apply the derivative to the outermost function, the entire square root thing:\[\Large\rm y'=\frac{1}{2\sqrt{3x+\sqrt{x^2+1}}}\]What I wrote is not correct yet, we have to apply the chain rule!!\[\Large\rm y'=\frac{1}{2\sqrt{3x+\sqrt{x^2+1}}}\color{royalblue}{\left(3x+\sqrt{x^2+1}\right)'}\]

zepdrix (zepdrix):

The blue part represents work we still need to do. We need to multiply by the derivative of the inner function.

OpenStudy (anonymous):

oh i see

zepdrix (zepdrix):

So how bout that blue part? Can you figure out the derivative of those two terms? We'll end up getting another chain it looks like

OpenStudy (anonymous):

give me one minute

OpenStudy (anonymous):

it would be:

OpenStudy (anonymous):

\[3+\frac{ x }{ \sqrt{x^2+1} }\] ???

zepdrix (zepdrix):

So the 2's canceled out when you did that? Ah yes looks nice!\[\Large\rm y'=\frac{1}{2\sqrt{3x+\sqrt{x^2+1}}}\color{royalblue}{\left(3+\frac{x}{\sqrt{x^2+1}}\right)}\]

zepdrix (zepdrix):

Good job! \c:/ You could probably do some extra work finding a common denominator and such, but that would be a nightmare. This is a good place to stop unless your teacher wants you to go further.

OpenStudy (anonymous):

My teacher wants it further :(

zepdrix (zepdrix):

Ahhh, weak sauce D:

zepdrix (zepdrix):

\[\Large\rm y'=\frac{3}{2\sqrt{3x+\sqrt{x^2+1}}}+\frac{x}{\sqrt{x^2+1}\cdot2\sqrt{3x+\sqrt{x^2+1}}}\]So I guess you can distribute... no no no, don't do that. that was silly of me.

zepdrix (zepdrix):

Let's get a common denominator in the blue part, combine, THEN distribute. much easier that way.

zepdrix (zepdrix):

\[\Large\rm y'=\frac{1}{2\sqrt{3x+\sqrt{x^2+1}}}\color{royalblue}{\left(\frac{3\sqrt{x^2+1}}{\sqrt{x^2+1}}+\frac{x}{\sqrt{x^2+1}}\right)}\] \[\Large\rm y'=\frac{1}{2\sqrt{3x+\sqrt{x^2+1}}}\color{royalblue}{\left(\frac{3\sqrt{x^2+1}+x}{\sqrt{x^2+1}}\right)}\]

zepdrix (zepdrix):

Something like that yah? and then just multiply across i guess?

zepdrix (zepdrix):

\[\Large\rm y'=\frac{3\sqrt{x^2+1}+x}{2\sqrt{x^2+1}\sqrt{3x+\sqrt{x^2+1}}}\]

zepdrix (zepdrix):

yah? :o make sense? Any confusing steps in there?

OpenStudy (anonymous):

i think I got it, but the teacher's answer was weird. \[\frac{ 3\sqrt{x^2+1}+x }{ 2\sqrt{3x(x^2+1)+(x^2+1)\sqrt{x^2+1}} }\]

zepdrix (zepdrix):

That seems so unnecessary of a step.. ugh. So teach combined the square roots. He/She took the (x^2+1) and distributed it to each term under the other root.

zepdrix (zepdrix):